OpenStudy (anonymous):

Help pleaseee algebra 1- ill give you a lot of medals and ill fan if you help me

OpenStudy (godlovesme):

ok i think graphing would be the easiest way to do this :D do u have geogebra app?

OpenStudy (godlovesme):

@blahhedup101 u there?

OpenStudy (anonymous):

yeah im here

OpenStudy (anonymous):

@Godlovesme

OpenStudy (anonymous):

and no i dont have a graphing app

OpenStudy (godlovesme):

hmm is it fine if i graph it and u determine how many solutions the line will have? solution is just the intersection so it's easy :)

OpenStudy (anonymous):

ok sounds good :)

OpenStudy (godlovesme):

great! :D so first are u familliar with the slope-intercept formula?

OpenStudy (godlovesme):

y= mx + b ?

OpenStudy (anonymous):

sorry my mom was talking to me

OpenStudy (anonymous):

but yes i am familiar with y=mx+b

OpenStudy (godlovesme):

it's cool :) so i guess you know that m stands for slope and b represents the y-intercept

OpenStudy (anonymous):

yes

OpenStudy (godlovesme):

awesome! so if two lines have the same slope, it means thy are parallel. And if two lines are parallel, they will NEVER intersect which means there will be no solution so do u see any lines with the same slope?

OpenStudy (anonymous):

oh ok that makes sense and idk i dont know how to graph them

OpenStudy (godlovesme):

no we dont need to graph it if we find any two equations that have the same slope for example y= 3x +1 and y= 3x +2 have the same slope,3 so the lines are parallel so there is no solution ( they will never intersect) make sense?

OpenStudy (anonymous):

ohhhhhhhhh

OpenStudy (anonymous):

ok ummm

OpenStudy (anonymous):

what about the last on is that no solution?

OpenStudy (godlovesme):

so if u look at the first pair, you can tell that they have the same slope because -x is the same thing as -1x

OpenStudy (anonymous):

and the second last one is no solution

OpenStudy (godlovesme):

hold on we're on the first one

OpenStudy (godlovesme):

y= -x -2 y= -x -6 do this equations have the same slope? ^^

OpenStudy (anonymous):

um no

OpenStudy (godlovesme):

um actually they do, look -x means -1x so we will have y= -1x -2 y= -1x -6 they both have a slope of -1

OpenStudy (anonymous):

ohhh... i thought they didnt because of the 2 and the 6

OpenStudy (godlovesme):

that's the y-intercept , the slope is the one with x :)

OpenStudy (anonymous):

i dont get these they are very confusing sooo the numbers dont count ?

OpenStudy (godlovesme):

sorry about the confusion :/ butt here y=mx+b we will be given numbers for m and b like y= 2x + 3 m = slope b= y-intercept so in this case our slope is 2 and the y-intercept is 3 got it?

OpenStudy (godlovesme):